### v1

@@ 7 Conclusions
A potential future direction is to develop Kaczmarz-type methods that can exploit not just large but also small isolated singular values to achieve fast convergence, as motivated by applications in partial differential equations, among others. Another natural question is whether it is possible to adapt Krylov subspace methods to take advantage of row-sampling techniques in order to improve their computational guarantees beyond what is possible through only matrix-vector products.
@@ Acknowledgments
Thanks to Daniel LeJeune for helpful discussions regarding accelerated sketch-and-project, and for sharing the programming environment. Also, thanks to Sachin Garg for helpful discussions on block memoization.
@@
where in the last step we use the fact that $\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}\|\leq 1/\mu$ and $\mu\geq\tilde{\mu}$. For the bound on ${\mathbf{e}}_{t}=\tilde{{\mathbf{w}}}_{t}-{\mathbf{w}}_{t}$, supposing that $\|{\mathbf{e}}_{t}\|\leq\epsilon_{1}\|{\mathbf{w}}_{t}\|$, we have
${\mathds{E}}[\|{\mathbf{e}}_{t}\|^{2}]\leq\epsilon_{1}^{2}{\mathds{E}}[\|{\mathbf{w}}_{t}\|^{2}]=\epsilon_{1}^{2}{\mathds{E}}[\|{\mathbf{P}}_{{\lambda},S}({\mathbf{x}}_{t}-{\mathbf{x}}^{*})\|^{2}]\leq\epsilon_{1}^{2}\|\bar{{\mathbf{P}}}_{{\lambda}}\|\cdot\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}$ (A.3)
@@
${\mathds{E}}[\|{\mathbf{e}}_{t}\|_{\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}}^{2}]\leq\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}\|\cdot{\mathds{E}}[\|{\mathbf{e}}_{t}\|^{2}]\leq\frac{\epsilon_{1}^{2}\|\bar{{\mathbf{P}}}_{{\lambda}}\|}{\tilde{\mu}}\cdot\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}$ (A.4)
@@
$~{}+(1+\phi)\underbrace{\left(\frac{1-\beta}{\tilde{\mu}}+\gamma^{2}\tilde{\nu}-2\gamma-\frac{\gamma\beta(1-\alpha)}{\alpha}+\frac{\gamma^{2}\epsilon_{1}^{2}\|\bar{{\mathbf{P}}}_{{\lambda}}\|}{\phi}(\tilde{\nu}+\frac{1}{\tilde{\mu}})\right)}_{P_{2}}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}.$
By choosing $\alpha=\frac{1}{1+\gamma\tilde{\nu}}$ and $\gamma=\frac{1}{\sqrt{\tilde{\nu}\tilde{\mu}}}$ we have $P_{1}=\gamma^{2}\tilde{\nu}=\frac{1}{\tilde{\mu}}$. By choosing $\beta=1-\sqrt{\frac{\tilde{\mu}}{\tilde{\nu}}}$ we have $\beta(1+\phi)\leq 1-\frac{1}{2}\sqrt{\frac{\tilde{\mu}}{\tilde{\nu}}}$ holds for $\phi=\frac{\sqrt{\tilde{\mu}/\tilde{\nu}}}{2(1-\sqrt{\tilde{\mu}/\tilde{\nu}})}$. By further setting $\epsilon_{1}\leq\frac{\tilde{\mu}}{4}\leq\frac{\tilde{\mu}}{\sqrt{8\|\bar{{\mathbf{P}}}_{{\lambda}}\|(\tilde{\mu}\tilde{\nu}+1)}}$ we also achieve $P_{2}\leq 0$. Finally, denote $\Delta_{t}=\|{\mathbf{v}}_{t}-{\mathbf{x}}^{*}\|_{\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}}^{2}+\frac{1}{\tilde{\mu}}\|{\mathbf{y}}_{t}-{\mathbf{x}}^{*}\|^{2}={\mathbf{r}}_{t}+\frac{1}{\tilde{\mu}}\|{\mathbf{y}}_{t}-{\mathbf{x}}^{*}\|^{2}$, we conclude that

### v2

@@ 7 Conclusions
A potential future direction is to develop Kaczmarz-type methods that can exploit not just large but also small isolated singular values to achieve fast convergence, as motivated by applications in partial differential equations, among others. Another natural question is whether it is possible to adapt Krylov subspace methods to take advantage of row-sampling techniques in order to improve their computational guarantees beyond what is possible through only matrix-vector products. Finally, large-scale implementation and experiments evaluating our methods in terms of wall-clock time, as well as for sparse problems, is an exciting next step.
@@ Acknowledgments
Thanks to Daniel LeJeune for helpful discussions regarding accelerated sketch-and-project, and for sharing the programming environment. Also, thanks to Sachin Garg for helpful discussions on block memoization. We also thank the editor and reviewers for their useful feedback that greatly improved the manuscript.
@@
where in the last step we use the fact that $\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}\|\leq 1/\mu$ and $\mu\geq\tilde{\mu}$. For the bound on ${\mathbf{e}}_{t}=\tilde{{\mathbf{w}}}_{t}-{\mathbf{w}}_{t}$, supposing that $\|{\mathbf{e}}_{t}\|\leq\epsilon_{1}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|$, we have
${\mathds{E}}[\|{\mathbf{e}}_{t}\|^{2}]\leq\epsilon_{1}^{2}\cdot{\mathds{E}}[\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}]=\epsilon_{1}^{2}\cdot\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}$ (A.3)
@@
${\mathds{E}}[\|{\mathbf{e}}_{t}\|_{\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}}^{2}]\leq\|\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}\|\cdot{\mathds{E}}[\|{\mathbf{e}}_{t}\|^{2}]\leq\frac{\epsilon_{1}^{2}}{\tilde{\mu}}\cdot\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}.$ (A.4)
@@
$~+(1+\phi)\underbrace{\left(\frac{1-\beta}{\tilde{\mu}}+\gamma^{2}\tilde{\nu}-2\gamma-\frac{\gamma\beta(1-\alpha)}{\alpha}+\frac{\gamma^{2}\epsilon_{1}^{2}}{\phi}(\tilde{\nu}+\frac{1}{\tilde{\mu}})\right)}_{P_{2}}\|{\mathbf{x}}_{t}-{\mathbf{x}}^{*}\|^{2}.$
By choosing $\alpha=\frac{1}{1+\gamma\tilde{\nu}}$ and $\gamma=\frac{1}{\sqrt{\tilde{\nu}\tilde{\mu}}}$ we have $P_{1}=\gamma^{2}\tilde{\nu}=\frac{1}{\tilde{\mu}}$. By choosing $\beta=1-\sqrt{\frac{\tilde{\mu}}{\tilde{\nu}}}$ we have $\beta(1+\phi)\leq 1-\frac{1}{2}\sqrt{\frac{\tilde{\mu}}{\tilde{\nu}}}$ holds for $\phi=\frac{\sqrt{\tilde{\mu}/\tilde{\nu}}}{2(1-\sqrt{\tilde{\mu}/\tilde{\nu}})}$. By further setting $\epsilon_{1}\leq\frac{\tilde{\mu}}{4}\leq\frac{\tilde{\mu}}{\sqrt{8(\tilde{\mu}\tilde{\nu}+1)}}$ we also achieve $P_{2}\leq 0$. Finally, denote $\Delta_{t}=\|{\mathbf{v}}_{t}-{\mathbf{x}}^{*}\|_{\bar{{\mathbf{P}}}_{{\lambda}}^{\dagger}}^{2}+\frac{1}{\tilde{\mu}}\|{\mathbf{y}}_{t}-{\mathbf{x}}^{*}\|^{2}={\mathbf{r}}_{t}+\frac{1}{\tilde{\mu}}\|{\mathbf{y}}_{t}-{\mathbf{x}}^{*}\|^{2}$, we conclude that